OpenStudy (anonymous):

describe the pattern in the numbers and write the next two numbers 0.0625 , 0.125 , 0.25 yes i know they myt be realli eaasy fr u guys bt im nt dat smart

OpenStudy (saifoo.khan):

Find the difference..

OpenStudy (anonymous):

im blank

OpenStudy (saifoo.khan):

Awesome. Head to bed. Sleep! :D

OpenStudy (anonymous):

nah i cant yo i been lazing off fr da past year lol

OpenStudy (saifoo.khan):

Past years suck! But they are useful before exams! :D

OpenStudy (anonymous):

no i dnt have examz i mean i kinda have to do everything frm home im a senior nd i have to do geomtry algebra two nd all that nd i dnt have much tym

OpenStudy (saifoo.khan):

Woah! That's easy as cake.

OpenStudy (anonymous):

maybe for you but im not smart i could do other things but math anthing in math is hard for me

OpenStudy (saifoo.khan):

I can understand the hardships though! ^_~

OpenStudy (anonymous):

lol :)

OpenStudy (saifoo.khan):

0.25^2 = 0.625

OpenStudy (saifoo.khan):

Hold on a sec.

OpenStudy (anonymous):

okay

OpenStudy (anonymous):

doubling each time i think

OpenStudy (anonymous):

it would be easier if you had these written as fraction \[\frac{1}{16},\frac{1}{8},\frac{1}{4}\]pattern would be clearer

OpenStudy (anonymous):

thats how it was on the sheet thankz alot :)

OpenStudy (saifoo.khan):

Might be Satellite cheated from that same sheet! :O

OpenStudy (anonymous):

lol nah i dnt think so these sheets r messed up cnt really cheat from

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

(:

OpenStudy (anonymous):

^_^

OpenStudy (saifoo.khan):

^_~

OpenStudy (anonymous):

*_*

OpenStudy (saifoo.khan):

Lol, Stop googling emotions!

OpenStudy (saifoo.khan):

^_~ <- Best.

OpenStudy (anonymous):

no im not i alredy kno these :p k

OpenStudy (anonymous):

hahaha yah dahs ma fav

OpenStudy (saifoo.khan):

Genius!

OpenStudy (saifoo.khan):

Lol.

OpenStudy (anonymous):

i kno in dis stuff lol

OpenStudy (saifoo.khan):

That's the reason i said, Genius Nisa.

OpenStudy (anonymous):

lol thank you thank you im honoured :D :p

OpenStudy (saifoo.khan):

Haha. Welcome.

OpenStudy (anonymous):

:)

OpenStudy (saifoo.khan):

(:

OpenStudy (saifoo.khan):

Off from here. C ya later sometime.. :)

OpenStudy (anonymous):

okay :) bye

OpenStudy (anonymous):

Okay so, 0.0625=625/10000 0.125=125/1000 0.25=25/100 Notice the changes: In the numerator: The number is DIVIDED by 5 EVERYTIME, how? 25x5=125 125x5=625? So the next two numbers will be: divide 25 by 5 = 5 And reduce 1 0 from 100, it is 10 So, 5/10 or 0.50 And finally, divide 5 by 5, it is 1 And reduce 1 0 from 10, it is 1. So the last two numbers are: 0.50, 1 FINAL ORDER: 0.0625, 0.125, 0.25,0.5,1 Hope this helps :) If not ask me over MSN I will help!

OpenStudy (anonymous):

thank u :)

OpenStudy (anonymous):

No problem in other words, just multiply each number by 10/5 or 2 to get the next!